Water at room temperature is pumped through a smooth pipe. The pipe has diameter of 10cm. The total length is10 km. The discharge is 80 m3 /h. A pump is installed at the inlet. At the exit of pump, the absolute pressure of the water is 25 atm. The exit exposes to the atmosphere and its elevation is 100 m higher than the inlet. Estimate the frictional head loss hf.
Answer #1

d = 10 cm = 0.1m

Q = A x V = 80 m3/h = 80/3600 m3/s = 0.02222 m3/s

A = (pi/4) d^2 = 3.142/4 * (0.1^2) = 0.007855 m2

V = Q/A = 0.02222/0.007855 = 2.8287 m/s

Reynolds number, Re = V x d/(viscosity of water)= 2.8287 x 0.1/ (8.9 x 10^-4) = 317. 83

Re <3000 which implies that it is laminar flow

Darcy- weisbach friction factor, f = 64/Re = 64/317.83 = 0.20136

Head loss due to friction , hf= f x (L/d) x (V^2/2g) = 0.20136 x (10000/0.1) x (2.8287^2)/2 x9.81 x 10^-3

hf= 8.211 m